Find implicit differentiation of 2y+cot(xy^2)=3xy

1
Expert's answer
2022-06-14T08:17:14-0400
"F(x,y)=2y+\cot(xy^2)-3xy"

"y'(x)=-\frac{F'_x}{F'_y}=-\frac{-\frac{y^2}{\sin^2(xy^2)}-3y}{2-\frac{2xy}{\sin^2(xy^2)}-3x}"

"y'(x)=\frac{y^2+3y\sin^2(xy^2)}{2\sin^2(xy^2)-2xy-3x\sin^2(xy^2)}"